OpenStudy (anonymous):

Determine whether the lines through the pairs of points are parallel, perpendicular, or neither. (1, 2) and (5, 4); (0, 6) and (2, 7)

OpenStudy (anonymous):

First, find the slope between each pair of points: Slope 1: (4-2)/(5-1) = 2/4 = 1/2 Slope 2: (7-6)/(2-0) = 1/2 Since the slopes are the same, this tells us that the lines are parallel.